OpenStudy (anonymous):

The population of deer in a certain national park can be approximated by the function p(x)=150(1.07)^x , where x is the number of years since 1995. In which year will the population reach 300? Hint: An answer such as 2002.4 would represent the year 2002.

OpenStudy (anonymous):

Put in 300 for p(x) 300=50(1.03)^x now just solve for x, know how to start?

OpenStudy (anonymous):

Kind of, but I in the equation wrong my bad, 300 = 150(1.07)^x Do we start off by dividing both sides by 150?

OpenStudy (anonymous):

Yep that would be a great start

OpenStudy (anonymous):

so that would leave me with 150 = (1.07)^x What would I do from there?

OpenStudy (anonymous):

I mean 2 = (1.07)^x

OpenStudy (anonymous):

Yep, now remember you can use logarithms to get that x out of the exponent by doing this: \[\log _{1.07}2=\log _{1.07}1.07^{x}\] the logarithm and base of exponent cancel on the right leaving \[\log _{1.07}2=x\]

OpenStudy (anonymous):

Oh yeah!Now do you write it in the exponential form?

OpenStudy (anonymous):

Well ya put that bad boy in a calculator and it gives you 10.2448 so 10 years since 1995=what year?

OpenStudy (anonymous):

2005?

OpenStudy (anonymous):

Thats right

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

You're very welcome